OpenStudy (anonymous):

Subject: systems with three variables For a party, you are cooking a large amount of stew that has meat, potatoes, and carrots. the meat costs $6 per pound. the potatoes cost $3 per pound and the carrots cost $1 per pound. you spend $48.50k on 13.5 pounds of food. you buy twice as many carrots as potatoes. ~ ALL I NEED ARE THE THREE EQUATIONS USED TO SOLVE.

OpenStudy (anonymous):

meat, m potatoes, p carrots, c $m = m*6 $p = p*3 $c = c*1 total cost = 48.50 = $m + $p + $c total pounds = 13.5 = m + p + c ya?

OpenStudy (anonymous):

thank you